OpenStudy (gabylovesyou):

Determine the type of boundary line and shading for the graph of the inequality y >/= 2x + 6 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm assuming that ">/= " is saying "greater than or equal to"?

OpenStudy (gabylovesyou):

yes :)

jimthompson5910 (jim_thompson5910):

so the boundary line is a solid one

jimthompson5910 (jim_thompson5910):

let's check the origin

jimthompson5910 (jim_thompson5910):

y >= 2x + 6 0 >= 2(0)+6 0 >= 0+6 0 >= 6 which is FALSE So the origin is NOT in the shaded region

OpenStudy (gabylovesyou):

B?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (gabylovesyou):

Thanks :)

jimthompson5910 (jim_thompson5910):

yw